# International school for advanced studies

### v1

@@ 7 Intrinsic dimension of hidden representations in deep neural networks
Using the tools described in the previous section, we study the ID of the ImageNet representations in a ResNet152, a standard architecture for image classification, and in Image GPT model (iGPT, [66]) an attention-based model trained to generate images. In many models for image classification, the ID grows in the first layers and then decreases, reaching its lowest value in the output layer [81]. In contrast, we will see that in the large iGPT architecture the profile of the ID has a "double hunchback" shape and that the two peaks arise in different stages during the training dynamics.
@@
In both cases, the ID decreases as the neighborhood size grows, more markedly for high ID values. At representation "34", when the ID reaches its peak, the scale analysis done with the TwoNN estimator is more unstable than that done with Gride. In panel 10-c we summarize the profile of ID across all the hidden representations choosing for simplicity, for all the representations, the values measured at the same scale: $k_{1}=2,k_{2}=4$ for Gride, $N=\nicefrac{{N_{tot}}}{{2}}$ for TwoNN.
@@
The ResNet152 is trained for 90 epochs on ImageNet with weight decay = $10^{-4}$, batch size = 256, and initial learning rate = 0.1. We decrease the learning rate by a factor of 10 at epoch $\{30,60,80\}$. The input size is $224\times 224$. We reach a top1 test accuracy of 78.22, consistent with Pytorch_ResNet152_v1. When the network is not trained, the ID after the first convolution decreases from 45 to 18. Then, after the first $8/13$ epochs the ID rapidly reaches its peak in all the layers, forming the hunchback shape observed at the end of training. The largest part of training is spent adjusting the ID of the representation toward the final value. This entails a steady decrease of the ID in the layers towards the output while in the first 10/30 layers the ID remains stable during training.
We conclude this section by showing the intra-class ID measured on the ImageNet training set in the space of the logits. The average ID is around 20 in the output space (see Fig. 10-c) but we do not know how much the ID differs from class to class. [81] showed that the ID of the output representation is correlated to the top5 test accuracy of the network: the lower the ID the higher the accuracy.
It is therefore interesting to test if this property holds also at the class level. Since the number of images per class in the validation set is only 50, we evaluate the IDs and accuracies on the training set where there are 895 classes that contain 1300 images and compute the ID with Gride using ($k_{1},k_{2})=(2,4)$. Figure 12-a shows that the intra-class IDs are quite variable with an average of 19,
@@
$p(x_{i})\in\mathbb{R}^{512}$ conditioned on the values of the previous pixel codes. The network is trained to predict pixel codes minimizing a negative log-likelihood loss [66]:
@@
We consider 90000 images (from 300 classes using 300 images per class) of the ImageNet training set, as we did for the ResNet architecture (see Sec. 1), and measure the ID of the representations after the pixel embedding, at the normalization layers before the attention blocks and at the output of the network. In all cases, we average over the sequence dimension, so that the size of the representation of an image is ($X_{l}\in\mathbb{R}^{D}$). Figure 13 shows the profile of the ID estimated with Gride $(k_{1},k_{2})=(2,4)$ for the small, medium, and large iGPT networks from left to right. In all cases, the IDs of the output are similar to those of the input.
Figure 13: Intrinsic dimension of Image GPT hidden layers ID evaluated on the normalization layers before some attention blocks of iGPT small (a), medium (b), large (c). The ID is measured with Gride setting $(k_{1}=2,k_{2}=4)$. As the size of the mode grows a second peak of the ID profile can be observed. The peak can be clearly detected for iGPT large around layers 41/44.
This is not surprising: since the network is trained to reconstruct a realistic completion of the images the geometric aspects of the data manifold like the ID should not be very different between output and input. Across the hidden layers the ID grows, but while on the small iGPT the ID profile has a hunchback shape quite similar to that observed in ResNet152 and in many convolutional discriminative models [81], on the medium and, most notably, large iGPT the ID has two peaks with a local minimum in the middle of the network. The formation of the two peaks occurs at different stages during training. Figure 14 shows the ID profile after $\{0,\>1.3,\>2.6,\>5.2,\>10\}\cdot 10^{5}$ iterations [^1] These checkpoints are available at https://github.com/openai/image-gpt. After $2.6\cdot 10^{5}$ iterations, the development of the first peak around layer 12 is completed but the second one at layer 40 is not yet formed. In the course of the last $7.4\cdot 10^{5}$ iterations, the intermediate minimum after block 23 slightly decreases from 25 to 24.2 and the ID at block 41 grows from 24.3 to 26.3 (iteration $5.2\cdot 10^{5}$) to finally reach 28.
@@
Figure 14: Evolution of the ID of iGPT representations during training. ID evaluated on the normalization layers before some attention blocks of iGPT Large during training. We measure the ID of the randomly initialized network, after $(1.3,2.6,5.2)\cdot 10^{6}$ iterations, and at the end of training. After $(2.6\cdot 10^{6}$ iterations, the first hunchback is formed and during the remaining training epochs, the second hunchback grows. The checkpoints are available at https://github.com/openai/image-gpt.
@@ 4 Probability landscape of other deep image models
In iGPT, the $ARI$ profiles are essentially zero in all the layers, meaning that, according to the class partition, the cluster assignments are random. Around layer 20 however, the overlap is 0.35 and the local composition of the neighborhoods is "ordered". This local ordering is, to some extent reflected by a homogeneous composition at the animals/objects level inside the middle-sized clusters: the large majority of clusters with sizes between 100 and 3000 have a homogeneity higher than 95% with the animals/objects partition.

### v2

@@ 7 Intrinsic dimension of hidden representations in deep neural networks
Using the tools described in the previous section, we study the ID of the ImageNet representations in a ResNet152, a standard architecture for image classification, and in Image GPT model (iGPT, [66]) an attention-based model trained to generate images. In many models for image classification, the ID grows in the first layers and then decreases, reaching its lowest value in the output layer [81]. In contrast, we will see that in the large iGPT architecture the profile of the ID has a "double hunchback" shape, and that the two peaks arise in different stages during the training dynamics.
@@
In both cases, the ID decreases as the neighborhood size grows, more markedly for high ID values. At representation "34", when the ID reaches its peak, the scale analysis done with the TwoNN estimator is more unstable than that done with Gride. In panel 10-c we summarize the profile of ID across all the hidden representations, choosing for simplicity, for all the representations, the values measured at the same scale: $k_{1}=2,k_{2}=4$ for Gride, $N=\nicefrac{{N_{tot}}}{{2}}$ for TwoNN.
@@
The ResNet152 is trained for 90 epochs on ImageNet with weight decay = $10^{-4}$, batch size = 256, and initial learning rate = 0.1. We decrease the learning rate by a factor of 10 at epochs $\{30,60,80\}$. The input size is $224\times 224$. We reach a top1 test accuracy of 78.22, consistent with Pytorch_ResNet152_v1. When the network is not trained, the ID after the first convolution decreases from 45 to 18. Then, after the first $8/13$ epochs the ID rapidly reaches its peak in all the layers, forming the hunchback shape observed at the end of training. The largest part of training is spent adjusting the ID of the representation toward the final value. This entails a steady decrease of the ID in the layers towards the output, while in the first 10/30 layers, the ID remains stable during training.
We conclude this section by showing the intra-class ID measured on the ImageNet training set in the space of the logits. The average ID is around 20 in the output space (see Fig. 10-c), but we do not know how much the ID differs from class to class. [81] showed that the ID of the output representation is correlated to the top5 test accuracy of the network: the lower the ID the higher the accuracy.
It is therefore interesting to test if this property holds also at the class level. Since the number of images per class in the validation set is only 50, we evaluate the IDs and accuracies on the training set where there are 895 classes that contain 1300 images, and compute the ID with Gride using ($k_{1},k_{2})=(2,4)$. Figure 12-a shows that the intra-class IDs are quite variable with an average of 19,
@@
$p(x_{i})\in\mathbb{R}^{512}$ conditioned on the values of the previous pixel codes. The network is trained to predict pixel codes by minimizing a negative log-likelihood loss [66]:
@@
We consider 90000 images (from 300 classes using 300 images per class) of the ImageNet training set, as we did for the ResNet architecture (see Sec. 1), and measure the ID of the representations after the pixel embedding, at the normalization layers before the attention blocks, and at the output of the network. In all cases, we average over the sequence dimension, so that the size of the representation of an image is ($X_{l}\in\mathbb{R}^{D}$). Figure 13 shows the profile of the ID estimated with Gride $(k_{1},k_{2})=(2,4)$ for the small, medium, and large iGPT networks from left to right. In all cases, the IDs of the output are similar to those of the input.
Figure 13: Intrinsic dimension of Image GPT hidden layers ID evaluated on the normalization layers before some attention blocks of iGPT small (a), medium (b), and large (c). The ID is measured with Gride setting $(k_{1}=2,k_{2}=4)$. As the size of the mode grows, a second peak of the ID profile can be observed. The peak can be clearly detected for iGPT large around layers 41/44.
This is not surprising: since the network is trained to reconstruct a realistic completion of the images the geometric aspects of the data manifold, like the ID, should not be very different between output and input. Across the hidden layers, the ID grows; but while on the small iGPT the ID profile has a hunchback shape quite similar to that observed in ResNet152 and in many convolutional discriminative models [81], on the medium and, most notably, large iGPT the ID has two peaks with a local minimum in the middle of the network. The formation of the two peaks occurs at different stages during training. Figure 14 shows the ID profile after $\{0,\>1.3,\>2.6,\>5.2,\>10\}\cdot 10^{5}$ iterations [^1] These checkpoints are available at https://github.com/openai/image-gpt. After $2.6\cdot 10^{5}$ iterations, the development of the first peak around layer 12 is completed but the second one at layer 40 is not yet formed. In the course of the last $7.4\cdot 10^{5}$ iterations, the intermediate minimum after block 23 slightly decreases from 25 to 24.2, and the ID at block 41 grows from 24.3 to 26.3 (iteration $5.2\cdot 10^{5}$) to finally reach 28.
@@
Figure 14: Evolution of the ID of iGPT representations during training. ID evaluated on the normalization layers before some attention blocks of iGPT Large during training. We measure the ID of the randomly initialized network, after $(1.3,2.6,5.2)\cdot 10^{6}$ iterations, and at the end of training. After $(2.6\cdot 10^{6}$ iterations, the first hunchback is formed, and during the remaining training epochs, the second hunchback grows. The checkpoints are available at https://github.com/openai/image-gpt.
@@ 4 Probability landscape of other deep image models
In iGPT, the $ARI$ profiles are essentially zero in all the layers, meaning that, according to the class partition, the cluster assignments are random. Around layer 20, however, the overlap is 0.35 and the local composition of the neighborhoods is "ordered". This local ordering is, to some extent reflected by a homogeneous composition at the animals/objects level inside the middle-sized clusters: the large majority of clusters with sizes between 100 and 3000 have a homogeneity higher than 95% with the animals/objects partition.